# Modelling Quantum Transduction for Multipartite Entanglement Distribution

### v1

@@ III-D TMD with intrinsic-entanglement quantum transduction & swapping (IES-TMD)
An additional model for TMD with intrinsic-entanglement quantum transduction is obtained by relaxing the assumption of concentrating intrinsic entanglement generation at the orchestrator, hence assuming that also clients are able to generate intrinsic EPR pairs between microwave and optical domains and assuming that additional quantum hardware is available along the optical links.
Accordingly, the generation of the EPR pairs occurs “at both points” rather than at “source only” [13, 47], as shown in Fig. 8. Specifically, $n$ intrinsic EPR pairs are generated at the orchestrator while $n$ intrinsic EPR pairs are generated at the clients as well. As for the IE-TMD scheme, each EPR pair is hybrid, involving both the microwave and optical domains. EPR pairs are eventually distributed between orchestrator and clients through a procedure resembling entanglement swapping [48]. This scheme, referred to as TMD with intrinsic-entanglement quantum transduction and swapping (IES-TMD), accounts for the nature of the Fock states in (8) [49], and it is described in details in the following.
@@
The optical ebits of the EPRs are thus transmitted through optical quantum channels and reach $n$ beam splitters, one for each client, with each beam splitter followed by two detectors. The overall setup is unable to distinguishing the which-path information [43, 49]. A click of one of the two detectors of each setup denotes the presence of an optical photon. However, due to the path-erasure – i.e., the impossibility of knowing whether the optical photon responsible for the detector-click has been generated at the orchestrator or at the client – it is impossible to distinguish where the SPDC has taken place (namely, whether at the orchestrator or at the client), and thus it is impossible to distinguish whether a microwave photon is present at orchestrator or at client. This results into the generation of another form of path-entanglement [50] between the microwave photons at the orchestrator and at the client [43]. Thus, the overall effect of beam splitter and detectors is reminiscent of entanglement swapping, since they project the received optical photons into a Bell state and the heralded signal – i.e., the detector-click – indicates the distribution of entangled pairs in the remote superconducting processors [51] in the form of:
|$\ket{\Psi_{MM}^{oc}}=\frac{1}{\sqrt{2}}\big{(}\ket{0_{M}^{o}1_{M}^{c}}+\ket{1_{M}^{o}0_{M}^{c}}\big{)}$ (13)
@@
It is important to highlight that, when both transducers generate optical photons, only one detector click is triggered due to path erasure. In such a case, a detector click corresponds to the presence of two microwave photons, one at the orchestrator and one at the client. Hence, the state shared between the orchestrator and the client is |$\ket{1^{o}_{M}}\ket{1^{c}_{M}}$, which is definitely not an entangled state as in (9). However, if we reasonably assume the availability of photon-counting detectors, then it is possible to distinguish the event of receiving two optical photons – one for each transducer in each link – from the event where only one optical photon is received. And the double-photon event can be discarded in favour of a new distribution attempt.
@@
$p^{IES}_{c}=S(\eta)*2\big{[}\eta-\eta^{2}\big{]}e^{-\frac{l_{o,c}}{2L_{o}}}$ (14)
where $l_{o,c}$ denotes the length of the fiber link between orchestrator and client $c$, $L_{o}$ denotes the attenuation length of the fiber, $\eta$ denotes the efficiency of the intrinsic entanglement generation at both orchestrator and clients, and $S(\cdot)$ denotes the Von Neuman entropy given in (11).
@@
The IES-TMD strategy allows a moderate reduction of the cooperativity $C$ requirements for enabling non-negligible entanglement distribution probability, but at the price of:
@@
Yet, it must be acknowledged that the “key” advantage of such a strategy lies in the possibility of heralding entanglement via off-the-shelf hardware – i.e., via photon counters – possibility that instead is not offered by the other protocols. Specifically, a detector click for each transduction attempt constitutes an indicator for identifying the generation of entanglement between orchestrator and client, without destroying it.
Furthermore, in Lemma 5 we considered the availability of single-counter detectors. If this hardware requirement cannot be satisfied, we can assume the availability of simpler single-photon detectors (SPD), which are not able to distinguish whether a click is due to one or two temporally-coincident photons. In this case, the probability of a single detector click includes also the even where both the transducers at orchestrator and client generate an optical photon, and therefore it can be expressed as follows:
@@
Both the photon-counter probability (see (28) in the appendices) and the photon-detector probability given in (15) are reported in Fig. 10 as a function of the cooperativity $C$. It is easy to note that the photon-detector probability is significantly larger than the photon-counter probability. Yet, in case of simple photon detectors, only a fraction $\frac{2\eta-2\eta^{2}}{2\eta-\eta^{2}}$ of clicks correspond to entanglement generation, with the remaining click fraction corresponding to a failed attempt. We can further generalize (15) by taking into account the detectors efficiency $\eta_{d}$, one of the main parameter to quantifying the performance of SPD [52]. Specifically, if we consider IDQ single photon detector with $\eta_{d}=0.25$ [53] running in gated [^10] So that we can reasonably neglect the dark-counts by synchronising the transduction attempt, i.e. the pump pulse, with the detection window. mode [52], we obtain the green curve in Fig. 11. Clearly, while both ideal- and realistic-photon detector click rates tends to $1$ as the cooperativity increases, the photon-counter click reaches its maximum equal to $\frac{1}{2}$ for $C=C_{th}$, as discussed above. Yet, the former click rates includes failed attempts that do not herald entanglement.
@@
We graphically summarise in Fig. 10 the performance of the different strategies in terms of achivable capacity, as a function of the cooperativity $C$. It is worthwhile to highlight again that the transducer efficiency does not provide sufficient granularity to grasp all the mechanisms/phenomena underlying a transducer process, as the type of encoding implemented within the transducer. Thus, we theoretically evaluated upper bound for the ebit distribution probabilities and thus for the quantum capacities. This the reason for which in the legend of Fig. 10 we indicated “capacity upper bounds”.
IV Conclusions
@@
Efficient microwave-optic quantum transduction is a key ingredient for the deployment of quantum networks, enabling the interconnection between remote superconducting nodes with optical channels. To this aim, in this paper we presented different communication models for multipartite entanglement distribution based on recent advances in quantum transduction devices to go beyond direct conversion. We have showed that TMD strategies shift from “preserving the multipartite state during frequency up- and down conversions” to “generating high-fidelity EPR pairs“ for teleporting the multipartite state. And, indeed, among the TMD strategies we analysed the IES-TMD, which exhibits the remarkable property of requiring $50\%$ conversion efficiency for achieving effective transduction. This results in a less stringent requirement on cooperativity, but at the cost of an entanglement probability that never reaches 1. Additionally, the exclusive advantage of IES-TMD scheme over all others is the ability to herald entanglement through detector clicks. Therefore, by switching from direct transduction to intrisic entanglement generation and performing entanglement swapping, we are able to address the stringent error probability constraints required by direct transduction. The conducted analysis is far from being exhaustive. The reason is that the hardware influences hugely the quantum transduction performance, and there is not a unique hardware solution available in literature. Nevertheless, we provided guidelines for the QT comparison from a communication engineering perspective. And we do hope that this paper will strike up a dialogue among the different research communities involved to converge on a standard reference model, which would be of paramount importance for quantum network development.
@@ Appendix A Proof of Lemma 2
By accounting for the specificity of the different functional blocks, it results that the cascade of up-conversion, fiber channel and down-conversion can be modeled as an overall quantum erasure channel (QEC), where the incoming ebit is replaced by the erasure state |$\ket{\epsilon}$ with probability $(1-p^{v}_{c})$. Formally, the output density matrix $\rho_{o}$ on a given orchestrator-client link is given by:
@@ Appendix C Proof of Lemma 4
To prove the Lemma we first need to prove that the efficiency $\eta^{o}_{\uparrow}$ in (5) governs the intrinsic entanglement generation at the orchestrator between microwave and optical domain. To this aim, we assume to generate intrinsic entanglement thought the transducer initialization with a microwave photon and its probabilistic up-conversion into an optical photon[43, 54], as discussed in the Remark of Sec.III-C (“red detuning”). This conversion leads to a beam splitter interaction Hamiltonian [^11] The interaction Hamiltonian is obtained in the condition of the rotating-wave approximation [27]. $H_{int}=g\sqrt{\langle n\rangle}(ab^{\dagger}+a^{\dagger}b)$, where $a$ and $b$ are the optical and microwave mode, respectively, and $a$ and $a^{\dagger}$ denote the lowering and raising operator. As done in Sec. III-B, we assume unitary extraction ratios $\zeta_{o}=\zeta_{m}=1$, i.e., null internal cavity losses for both optical and microwave fields. To properly analyse the transducer, we exploit the input-output relations in the Heisenberg-Langevin form [27]:
@@
with $d=\frac{\gamma_{o}\gamma_{m}}{4}+g^{2}\langle n\rangle$. By accounting for the expression of $\eta^{o}_{\uparrow}$ in eq. (1), it results that $|L_{o,m}|=|L_{m,o}|=\sqrt{\eta^{o}_{\uparrow}}$ for any value of the extraction ratios $\zeta_{x}$, whereas $|L_{o,o}|=|L_{m,m}|$ follows from our assumption [^13] In general, when $\zeta_{x}<1$, $|L_{o,o}|\neq|L_{m,m}|$, due to the presence in (22) of the parameters $\gamma_{x,e}$. of unitary extraction ratios.
@@
and, in this condition the transducer acts as a 50/50 lossless beam-splitter [59] generating a Bell State |$\ket{\psi_{out}}=\frac{1}{\sqrt{2}}(\ket{0^{o}_{M}1^{o}_{O}}+\ket{1^{o}_{M}0^{o}_{O}})$ that it is equivalent to (8), unless a basis change. From (24), for any setting of operative parameters resulting in $\eta^{o}_{\uparrow}=\frac{1}{2}$ – i.e., $C=C_{th}=3-2\sqrt{2}$ – an EPR is generated. Conversely, whenever $C\neq C_{th}$, we have two cases. For degenerate values of $\eta^{o}_{\uparrow}$ – i.e., $\eta^{o}_{\uparrow}=0$ or $1$ – the transducer outputs a single mode and no entanglement is generated, whereas for non-degenerated values of $\eta^{o}_{\uparrow}$ an odd superposition of states |$\ket{01}$ and |$\ket{10}$ – which is not a maximally entangled state – is generated.
For quantify the entanglement generation probability, we resort to the concept of entanglement of distillation [60], which is an upper bound of the number $n$ of EPRs that can be distilled from $m$ copies of a pure state using LOCC. Formally, for a pure state with density matrix $\rho$, the distillable entanglement $E(\rho)$ is given by [61]:
@@
and thus (26) coincides with (11). The overall proof follows by accounting for (26), by regarding $\eta^{c}_{\downarrow}$ as the probability of converting an input optical pho ton into an output microwave photon at the client and by accounting for the length of the fiber between the orchastrator and the client.

### v2

@@ III-D TMD with intrinsic-entanglement quantum transduction & swapping (IES-TMD)
An additional model for TMD with intrinsic-entanglement quantum transduction is obtained by relaxing the assumption of concentrating intrinsic entanglement generation at the orchestrator, hence assuming that also clients are able to generate intrinsic EPR pairs between microwave and optical domains, and assuming that additional quantum hardware is available along the optical links.
Accordingly, the generation of the EPR pairs occurs “at both points” rather than at “source only” [13, 62], as shown in Fig. 8. Specifically, $n$ intrinsic EPR pairs are generated at the orchestrator while $n$ intrinsic EPR pairs are generated at the clients as well. As for the IE-TMD scheme, each EPR pair is hybrid, involving both the microwave and optical domains. EPR pairs are eventually distributed between orchestrator and clients through a procedure resembling entanglement swapping [63]. This scheme, referred to as TMD with intrinsic-entanglement quantum transduction and swapping (IES-TMD), accounts for the nature of the Fock states in (9) [64], and it is described in details in the following  [^12] As mentioned in the Remark of Sec. III-C, as a seek of clarity we consider the beam splitter interaction for the hybrin EPR pair generation within the transducer hardware..
@@
The optical ebits of the EPR pairs are thus transmitted through optical quantum channels and reach $n$ beam splitters, one for each client, with each beam splitter followed by two detectors. The overall setup is unable to distinguishing the which-path information [57, 64]. A click of one of the two detectors of each setup denotes the presence of an optical photon. However, due to the path-erasure – i.e., the impossibility of knowing whether the optical photon responsible for the detector-click has been generated at the orchestrator or at the client – it is impossible to distinguish where the up-conversion process has been successful (namely, whether at the orchestrator or at the client), and thus it is impossible to distinguish whether a microwave photon is present at orchestrator or at client. This results into the generation of another form of path-entanglement [65] between the microwave photons at the orchestrator and at the client [57]. Thus, the overall effect of beam splitter and detectors is reminiscent of entanglement swapping, since they project the received optical photons into a Bell state and the heralded signal – i.e., the detector-click – indicates the distribution of entangled pairs in the remote superconducting processors [66] in the form of:
|${\ket{\Psi_{M,M}^{o,c}}=\frac{1}{\sqrt{2}}\big{(}\ket{0_{M}^{o}1_{M}^{c}}+\ket{1_{M}^{o}0_{M}^{c}}\big{)}}$ (14)
@@
It is important to highlight that, when both transducers generate optical photons, only one detector click is triggered due to path erasure. However, in this case, the state shared between the orchestrator and the client is |$\ket{0^{o}_{M}}\ket{0^{c}_{M}}$, which is definitely not an entangled state as in (14). However, if we reasonably assume the availability of photon-number-resolved detectors (PNRD), then it is possible to distinguish the event of receiving two optical photons – one for each transducer in each link – from the event where only one optical photon is received. And the double-photon event can be discarded in favour of a new distribution attempt.
@@
$p^{IES}_{c}=S(\eta)*2\big{[}\eta-\eta^{2}\big{]}e^{-\frac{l_{o,c}}{2L_{0}}}$ (15)
where $l_{o,c}$ denotes the length of the fiber link between orchestrator and client $c$, $L_{0}$ denotes the attenuation length of the fiber, $\eta$ denotes the efficiency of the intrinsic entanglement generation at both orchestrator and clients, and $S(\cdot)$ denotes the Von Neuman entropy given in (12).
@@
TABLE II: EPR pairs distribution strategies from orchestrator to clients in different TMD strategies.
@@
In summary the IES-TMD strategy allows a moderate reduction of the cooperativity $C$ requirements for enabling non-negligible entanglement distribution probability, but at the price of:
@@
Yet, it must be acknowledged that the “key” advantage of such a strategy lies in the possibility of heralding entanglement via off-the-shelf hardware – i.e., via PNRD. Specifically, a detector click for each transduction attempt constitutes an indicator for identifying the generation of entanglement between orchestrator and client, without destroying it.
Also in Vanvilla-TMD and IE-TMD, it is possible to herald the entanglement with photon detectors by exploiting other degree of freedom such as time-bin [67, 42]. This also implies the introduction of additional heralding setups. On the contrary, in IES-TMD strategy entanglement heralding is embedded in the setup itself and, for this reason, it is not necessary to exploit other degree of freedom and/or to introduce any additional heralding equipment.
Furthermore, it is worth emphasizing that, in Lemma 5 we considered the availability of PNRD. If this hardware requirement cannot be satisfied, we can assume the availability of simpler single-photon detectors (SPD), which are not able to distinguish whether a click is due to one or two temporally-coincident photons. In this case, the probability of a single detector click includes also the event where both the transducers at orchestrator and client generate an optical photon, and therefore it can be expressed as follows:
@@
Both the PNRD probability (see (29) in the Appendix D) and the SPD probability given in (16) are reported in Fig. 10 as a function of the cooperativity $C$. It is easy to note that the SPD probability is significantly larger than the PNRD probability. Yet, in case of single photon detectors, only a fraction $\frac{2\eta-2\eta^{2}}{2\eta-\eta^{2}}$ of clicks correspond to entanglement generation, with the remaining click fraction corresponding to a failed attempt. We can further generalize (16) by taking into account the detectors efficiency $\eta_{d}$, one of the main parameter to quantifying the performance of SPD [68]. Specifically, if we consider IDQ single photon detector with detection efficiency $\eta_{d}=0.25$ [69] running in gated [^13] So that we can reasonably neglect the dark-counts by synchronising the transduction attempt, i.e. the pump pulse, with the detection window. mode [68], we obtain the green curve in Fig. 10. Clearly, while both ideal- and realistic-SPD click rates tends to $1$ as the cooperativity increases, the PNRD click reaches its maximum equal to $\frac{1}{2}$ for $C=C_{th}$, as discussed above. Yet, the former click rates includes failed attempts that do not herald entanglement.
TABLE III: Strategies comparison in terms of achievable quantum channel capacity and the corresponding required cooperativity $C$.
IV Discussion
Stemming from the developed analysis, it appears that the nature of the entangled pairs exploited in the TMD strategies are diverse. More into details, in the case of Vanilla-TMD, the EPR pairs, generated by the orchestrator, are at microwave frequencies. Therefore, in order to be distributed to the clients, one ebit for each EPR pair has to be up-converted to optical frequency through the transducer available at the orchestrator. On the contrary, in the case of IE-TMD and IES-TMD strategies, hybrid entangled pairs are generated: one ebit of each EPR pair is at microwave frequency and the other ebit is at optical frequency. Thus, by adopting IE-TMD or IES-TMD strategy, the EPR generation process is directly performed at the transducer hardware, by exploiting two mode squeezing or beam splitter interaction, respectively, as detailed in Sec. III-C. The aforementioned difference in the type and nature of the exploited EPR pairs in the different TMD strategies are summarized in Tab.II.
We graphically summarise in Fig. 11 the communication performances of the proposed strategies, in terms of achievable quantum capacity, as a function of the cooperativity $C$.
@@ IV Discussion
It is worthwhile to stress that the transducer efficiency does not provide sufficient granularity to grasp all the mechanisms/phenomena underlying a transducer process, as the type of encoding implemented within the transducer. Thus, we theoretically evaluated upper bound for the ebit distribution probabilities and thus for the quantum capacities. This is the reason for which in the legend of Fig. 11 we indicated “capacity upper bounds”. For the sake of clarity, Tab. III shows the strategies comparison in terms of achievable quantum capacity for the required values of cooperativity $C$.
The aforementioned consideration allows us to further stress an aspect highlighted in the final remark of Sec. III-A. Specifically, as observed, quantum transduction is not just a merely frequency conversion process. Indeed, several factors impact the transducer conversion efficiency and consequently the failure conversion probability [11, 70]. Nevertheless, the hardware literature still lacks of a general model able to grasp all the mechanisms/phenomena underlying the electro-optical transducer process from the entanglement generation and distribution perspective. Stemming from this, and by considering that our main goal is to start a dialogue among different communities (i.e. communication, physics, and hardware), we choose to propose a first model of quantum transducer in which the conversion efficiency constitutes the main parameter for characterizing its performances. The rational for this choice is that several factors determining the conversion failure can be embedded within the conversion efficiency. An example to all, to ensure a successful conversion, the mode phase-matching condition has to be fulfilled. The effect of mode mismatching is typically mitigated through filtering techniques [71], that reduces the probability of conversion and consequently the conversion efficiency. As result, the mode-mismatch can be embedded in the conversion efficiency parameter through the parameter $g$ in (1), as shown in [11]. Our design choice allows not only to abstract from the particulars of the specific technology underlying the transducer hardware, but also to easily and progressively take into account technological advancements that improve the transducer performances.
V Conclusions
@@
Efficient microwave-optic quantum transduction is a key ingredient for the deployment of quantum networks, enabling the interconnection between remote superconducting nodes with optical channels. To this aim, in this paper we presented different communication models for multipartite entanglement distribution based on recent advances in quantum transduction devices to go beyond direct conversion. We have showed that TMD strategies shift from “preserving the multipartite state during frequency up- and down conversions” to “generating high-fidelity EPR pairs" for teleporting the multipartite state. And, indeed, among the TMD strategies we analysed the IES-TMD, which exhibits the remarkable property of requiring $50\%$ conversion efficiency for achieving effective transduction. This results in a less stringent requirement on cooperativity, but at the cost of an entanglement probability that never reaches 1. Additionally, the exclusive advantage of IES-TMD scheme over all others is the ability to herald entanglement through detector clicks. Therefore, by switching from direct transduction to intrisic entanglement generation and performing entanglement swapping, we are able to address the stringent error probability constraints required by direct transduction. The conducted analysis is far from being exhaustive. The reason is that the hardware influences hugely the quantum transduction performance, and there is not a unique hardware solution available in literature. Nevertheless, we provided guidelines for the QT comparison from a communication engineering perspective. And we do hope that this paper will strike up a dialogue among the different research communities involved to converge on a standard reference model, which would be of paramount importance for quantum network development.
@@ Appendix A Proof of Lemma 2
By accounting for the specificity of the different functional blocks, it results that the cascade of up-conversion, fiber channel and down-conversion can be modeled as an overall quantum erasure channel
(QEC), where the incoming ebit is replaced by the erasure state |$\ket{\epsilon}$ with probability $(1-p^{v}_{c})$. Formally, the output density matrix $\rho_{o}$ on a given orchestrator-client link is given by:
@@ Appendix C Proof of Lemma 4
To prove the Lemma we first need to prove that the efficiency $\eta^{o}_{\uparrow}$ in (3) governs the intrinsic entanglement generation at the orchestrator between microwave and optical domain. To this aim, we assume to generate intrinsic entanglement thought the transducer initialization with a microwave photon and its probabilistic up-conversion into an optical photon [57, 72], as discussed in the Remark of Sec.III-C (“red detuning”). This conversion leads to a beam splitter interaction Hamiltonian [^14] The interaction Hamiltonian is obtained in the condition of the rotating-wave approximation [33]. $H_{int}=g\sqrt{\langle n\rangle}(ab^{\dagger}+a^{\dagger}b)$, where $a$ and $b$ are the optical and microwave mode, respectively, and $a$ and $a^{\dagger}$ denote the lowering and raising operator. As done in Sec. III-A, we assume unitary extraction ratios $\zeta_{o}=\zeta_{m}=1$, i.e., null internal cavity losses for both optical and microwave fields. To properly analyse the transducer, we exploit the input-output relations in the Heisenberg-Langevin form [33]:
@@
with $d=\frac{\gamma_{o}\gamma_{m}}{4}+g^{2}\langle n\rangle$. By accounting for the expression of $\eta^{o}_{\uparrow}$ in  (1), it results that $|L_{o,m}|=|L_{m,o}|=\sqrt{\eta^{o}_{\uparrow}}$ for any value of the extraction ratios $\zeta_{x}$, whereas $|L_{o,o}|=|L_{m,m}|$ follows from our assumption [^16] In general, when $\zeta_{x}<1$, $|L_{o,o}|\neq|L_{m,m}|$, due to the presence in (23) of the parameters $\gamma_{x,e}$. of unitary extraction ratios.
@@
and, in this condition the transducer acts as a 50/50 lossless beam-splitter [77] generating a Bell State |$\ket{\psi_{out}}=\frac{1}{\sqrt{2}}(\ket{0^{o}_{M}1^{o}_{O}}+\ket{1^{o}_{M}0^{o}_{O}})$. From (25), for any setting of operative parameters resulting in $\eta^{o}_{\uparrow}=\frac{1}{2}$ – i.e., $C=C_{th}=3-2\sqrt{2}$ – an EPR pair is generated. Conversely, whenever $C\neq C_{th}$, we have two cases. For degenerate values of $\eta^{o}_{\uparrow}$ – i.e., $\eta^{o}_{\uparrow}=0$ or $1$ – the transducer outputs a single mode and no entanglement is generated, whereas for non-degenerated values of $\eta^{o}_{\uparrow}$ an odd superposition of states |$\ket{01}$ and |$\ket{10}$ – which is not a maximally entangled state – is generated.
For quantify the entanglement generation probability, we resort to the concept of entanglement of distillation [78], which is an upper bound of the number $n$ of EPR pairs that can be distilled from $m$ copies of a pure state using LOCC. Formally, for a pure state with density matrix $\rho$, the distillable entanglement $E(\rho)$ is given by [79]:
@@
and thus (27) coincides with (12). The overall proof follows by accounting for (27), by regarding $\eta^{c}_{\downarrow}$ as the probability of converting an input optical photon into an output microwave photon at the client and by accounting for the length of the fiber between the orchestrator and the client.